Telescope and AGw Commissioning at LFBG with IRTC - 13-14 April 2009 MST

Observers: D.Thompson (@LBT)
Telescope Operator: W.Wack
Telescope Support:
Instrument Support: JMorris
SW Support: TLeibold

Summary

  • Lost most of the night to heavy clouds.
  • Another set of data on M67 shows the same high rms in X.

Details

Calibration unit has been removed.

18:50 MST (Sunset) Opening dome to patchy but thick cirrus that looks to be with us for the whole night. Tair=3.6 and falling. Tmirror=3.5 RH=41%, Wind=5m/s avg. Vent doors are open as well.

18:57 Right shutter overcurrent problem, plus Waytne and John are going to check that the bolts from the IRTC to the AGW are tight. One bolt on the AGW-to-Rotator found to be loose (backed off ~1/4-inch, Wayne tightened it by hand but this should be checked tomorrow - see IT# ).

19:35 ACQUIRE mode preset sent to M67astrom field. Failed (M1 failed because chase state not IDLE). My fault, I sent the preset too early.

19:38 ACQUIRE mode preset sent again, same source. Adjust M1 focus -0.5mm on global offset. There is a second source nearby, so...

19:42 ACQUIRE mode preset to WT10_254. Nothing there, spiral search

19:51 TRACK mode preset to M5_0347 to figure out pointing. May be cloudy. Star acquired at -2,+15 arcsec RADEC. IE adjusted from -70 to -85, CA from -69 to -49

19:58 pointto WT10_254 in ACQUIRE mode. IE now -78 CA now -46

20:00 pointto M67astrom gs=35 ACTIVE - successful. Collimating: 3371nm (mostly focus), 1680, 905, 414,

20:38 Re-correct pointing before starting AGW calibration run. IE now -83, CA now -49.5

20:40 PA=0 on M67astrom 0.5x40 datacubes EL~63 at start, LST=10:00 so the field is setting. 0.75 arcsec on guider. Verified that the field stars have proper motions in the catalog and that they are being applied.

GS cube GP_X GP_Y 35 0004 17.074 572.923
77 0005 117.667 549.121
52 0006 44.034 489.911 picked wrong guide star (brighter star to SE)
46 0007 26.321 472.447
18 0008 -81.461 478.385
05 0009 -143.184 587.064
24 0010 -71.976 597.473
35 0011 -17.005 572.920
49 0012 30.452 450.008
78 0013 130.164 482.479
63 0014 80.076 554.470
47 0015 27.795 513.531
35 0016 -17.046 572.914
50 0017 40.717 585.448
27 0018 -47.644 590.857

Preset appears hung, these are the last messages in GCS syslog:
Apr 13 21:28:16 lbtmu107 LBT_left_GCS: checking _azcam member
Apr 13 21:28:16 lbtmu107 LBT_left_GCS: this _azcam instance is waiting for an image. _azcam deletion delayed by 32716 ms ...
Apr 13 21:28:16 lbtmu107 LBT_left_GCS: waiting for image to arrive...
Called Torsten. Image seems to not be getting a timeout. Restarted GCS (~21:36).

22 0019 -73.642 565.687
20 0020 -74.345 543.685
21 0021 -72.539 516.541
08 0022 -122.905 496.061

Need to correct pointing...IE now -90.5 , CA now -58.3.

Preset appears hung again, these are the last messages in GCS syslog:
Apr 13 21:54:55 lbtmu107 LBT_left_GCS: setting up imageserver IP
Apr 13 21:54:55 lbtmu107 LBT_left_GCS: init and connect ...
Apr 13 21:54:55 lbtmu107 LBT_left_GCS: Invalid handle to blockstub: )
Apr 13 21:54:55 lbtmu107 LBT_left_GCS: Invalid handle to blockstub: )
Restarted GCS again (21:59).

08 0022 -122.905 496.061
10 0023 -108.097 484.038
35 0024 -16.936 572.898
32 0025 -28.555 550.034
31 0026 -28.694 502.461
36 0027 -11.242 478.994

Note: the IRTC looks a little out of focus. Using movie mode, AIP_IRTC_L.cfg focus term changed from 35.35 to 35.31

68 0028 97.285 545.066
80 0029 133.279 507.150
74 0030 115.054 511.327
48 0031 28.574 536.739
35 0032 -16.921 572.873
35 0033 -16.907 572.870
35 0034 -16.866 572.868
35 0035 -16.866 572.864
35 0036 -16.866 572.858

Now authoriuze as LUCIFER and repeat a subset: IE and CA set back to the same values as for IRTC (-90.5, -58.3) as a first guess. Nope IE=-75.5 CA=12.3. Corrected at M67 to IE=-69.5, CA=8.0. Allow to collimate on axis

23:30 Nope...it is completely overcast now. M5 star at 61d EL is faint in 0.5s on IRTC.

-- 0037 sky for above, taken against thick clouds.

03:48 Reopened for some short tests through a "sucker hole"

Preset to pointing star - EL got stuck at 70 degrees. Stopped and restarted GCS only to get out of the hanging preset in iraf. Stuck again at 72 degrees EL. Trying a different star. Stuck again at EL~74.

Spiral search past 50" failed to find 7.5m star...clouds. Only a few stars are visible outside.

M5_1459 to the south...OK

04:40 We manage to get collimated on BS9170 to the south, going back to M92astrom. Way out of focus (half the acquisition image field of view). We were collimated (~240nm rms) at the previous position with 0.6 arcsec FWHM. Why was there such a huge jump in focus and coma?

Re-determine rotator center on IRTC, measured at EL~72 cube 39 PA=270 cube 40 PA=90 Only 2 pairs of stars...rotator center works out to 155.5 136.5.

05:18:21.374 2009 Preset command started for left LUCIFER mode TRACK from IIF
Track mode preset failed to return complete to the instrument. And the GCS syslog shows:
Apr 14 05:18:27 lbtmu107 LBT_left_GCS: IST: connect accepted!
Apr 14 05:18:27 lbtmu107 LBT_left_GCS: checking _azcam member
Apr 14 05:18:27 lbtmu107 LBT_left_GCS: this _azcam instance is waiting for an image. _azcam deletion delayed by 32716 ms ...
Apr 14 05:18:27 lbtmu107 LBT_left_GCS: waiting for image to arrive...
Why was GCS trying to take an image 6 seconds after the TRACK mode preset was sent?
GCS stopped and restarted, this did not clear the hung preset.

05:30 Closing.

-- DavidThompson - 12 Apr 2009
Topic revision: r5 - 14 Apr 2009, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback